OpenStudy (anonymous):

find the slope of the line through each pair of points (Please show your work!) (8,10), (-7,14) (-3,1), (-17,2) (-20,-4) , (-12, -10) (-12, -5), (0,-8) (-19, -6), (15,16) (-6,9), (7,-9) (-18, -20), (-18, 15) (12, -18), (11,12)

OpenStudy (anonymous):

this a lot of problems

myininaya (myininaya):

oh mine

OpenStudy (anonymous):

\[\frac{14-10}{-7-8}=\frac{4}{-15}\]

OpenStudy (anonymous):

there's one

myininaya (myininaya):

faster satellite!

OpenStudy (anonymous):

ok fine i do the odd ones.

myininaya (myininaya):

\[\frac{2-1}{-17-(-3)}=\frac{1}{-14}\]

OpenStudy (anonymous):

\[\frac{-10+4}{-12+20}=\frac{-6}{8}=\frac{-3}{4}\]\]

myininaya (myininaya):

\[\frac{-5-(-8)}{-12-0}=\frac{3}{-12}=\frac{-1}{4}\]

OpenStudy (anonymous):

\[\frac{16+6}{15+19}=\frac{22}{34}=\frac{11}{17}\]

myininaya (myininaya):

\[\frac{-9-9}{7-(-6)}=\frac{-18}{13}\]

OpenStudy (anonymous):

Undefine! done

myininaya (myininaya):

\[\frac{-18-12}{11-12}=\frac{-20}{-1}=20\]

myininaya (myininaya):

last one should be -20 since the bottom was suppose to be 12-11

OpenStudy (anonymous):

Which ones were those? i dont see the problem so im not sure which is which x.x

myininaya (myininaya):

just look for where you see the two pairs for each problem they should be in order

OpenStudy (anonymous):

i did the odd ones, 1,3,5,7

OpenStudy (anonymous):

Wha.

OpenStudy (anonymous):

OHHHH now i see it lol

OpenStudy (anonymous):

ty lmao